Information on Gerken Power Co., is shown below. Assume the company's tax rate is 38%.

Debt: 8,800 8.1% coupon bonds outstanding, $1,000 par value, selling for 103.5% of par, and YTM OF 7.7%.

Common Stock: 213,000 shares outstanding, selling for $83.30 per share; beta is 1.18.

Preferred Stock: 12,300 shares of 5.9% preferred stock outstanding, currently selling for $97.70 per share.

Market: 7.15% market risk premium and 4.95% risk-free rate.

A. What is the market value of the firm?

B. What is the company's cost of each form of financing?

C. Calculate the company's WACC?
Answer #1

A

MV of equity=Price of equity*number of shares outstanding
MV of equity=83.3*213000
=17742900
MV of Bond=Par value*bonds outstanding*%age of par
MV of Bond=1000*8800*1.035
=9108000
MV of Preferred equity=Price*number of shares outstanding
MV of Preferred equity=97.7*12300
=1201710
MV of firm = MV of Equity + MV of Bond+ MV of Preferred equity
=17742900+9108000+1201710
=28052610

B

Cost of equity
As per CAPM
Cost of equity = risk-free rate + beta * (Market risk premium)
Cost of equity% = 4.95 + 1.18 * (7.15)
Cost of equity% = 13.39
After tax cost of debt = cost of debt*(1-tax rate)
After tax cost of debt = 7.7*(1-0.38)
= 4.774
cost of preferred equity
cost of preferred equity = Preferred dividend/price*100
cost of preferred equity = 5.9/97.7*100
=6.04

C

Weight of equity = MV of Equity/MV of firm
Weight of equity = 17742900/28052610
W(E)=0.6325
Weight of debt = MV of Bond/MV of firm
Weight of debt = 9108000/28052610
W(D)=0.3247
Weight of preferred equity = MV of preferred equity/MV of firm
Weight of preferred equity = 1201710/28052610
W(PE)=0.0428
WACC=after tax cost of debt*W(D)+cost of equity*W(E)+Cost of preferred equity*W(PE)
WACC=4.77*0.3247+13.39*0.6325+6.04*0.0428
WACC =10.28%